OpenStudy (anonymous):

Juan is 5 years older than his sister. Ten years ago, he was twice as old as his sister. Find his present age.

OpenStudy (anonymous):

20

OpenStudy (anonymous):

These ones really hurt my head

OpenStudy (anonymous):

suppose the sister's age is S years. how old was she 10 years ago? If she is S years old, then Juan is S+5 years old.

OpenStudy (anonymous):

I still dont get it :(

OpenStudy (anonymous):

are we assuming she started at 0?

OpenStudy (anonymous):

I work it in a weird order. 5 years apart, then I multiply by 2, so that gave me that he was 10 when she was 5. Then I just add the 10 years since it said that was 10 years ago. So, 5 * 2 = 10 + 10 = 20

OpenStudy (anonymous):

OOO that makes sense, only when you told me it though..thanks for the help :)

OpenStudy (anonymous):

Start with this: If they're asking for two unknowns, it means you have to create two equations. Now look at the text of the problem. They're giving you mathematical data disguised as just words. "Juan is 5 years older than his sister" You can rewrite this as the equation: J=S+5 It's the same thing, we're just using J=Juan and S=Sister "Ten years ago, he was twice as old as his sister. J-10=2*(S-10) Now, to solve, since you know J=S+5 it means you can rewrite the second equation as: J-10=2*(S-10) Substituting the value for J from the first equation S+5-10=2S-20 Moving the S to the right and the numbers to the left... 5-10+20=2S-S 15=S Current age of Juan's sister is 15 Since you know J=S+5 it means J=15+5 =20 And if you subtract 10 from each, J(ten years ago)=10 and S(ten years ago)=5

OpenStudy (anonymous):

yay I was on the right path I just got stuck on j-10 2(s+5) I put 5 instead of 10

OpenStudy (anonymous):

thank you for your help